

********************************************************************
UARS WEEKLY STATUS REPORT - 16 JUL 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE began day only operations on Thursday in support of the yaw
maneuver. This mode continues until July 21. Otherwise, the instrument
continues to perform exceptionally, taking solar data on all available
orbital tracking periods. An average of 63% of data were returned this
week accounting for 495 minutes of solar observing time.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2003289. MEPS data files between
2003290 and 2003365 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 03 July 2004 (Orbit 70096, GMT Day 185) through 09
July 2004 (Orbit 70202, GMT Day 191).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 186/1631 70122 POWER ON / BORESITE GAINS
The beta angle ranged between the angles of 63.8 degrees to 36.5
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 18 July 2004 (DOY 200).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.27 to 1.28 degrees C. Additional
battery performance data for each battery on 02 July and 09 July 2004
are provided in Appendix D.
The clock error ranged from 10.1 to –11.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
• OUT-OF-LIMITS = 00
• ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
• FOT = 04
• HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
• FOT = 05
• SOLSTICE = 01
Four (4) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Six (6) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The database on the new Eclipse software build was validated last week
for both command and telemetry. With this step completed, the FOT
will begin testing the DRs fixed in this build. The goal for the
testing to be complete is the end of the July. At that time this new
build will be released to operations.
MEETINGS
The FOT attended the UARS Monthly Status Review on 05 July 2004 to
discuss the significant events for the month of June.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 July 2004 (DOY 199). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 09 July 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
70099 185/0254 04-198 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
70101 185/0656 04-199 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
70102 186/1631 04-200 HALOE POWER ON / BORESITE GAINS
70134 187/1149 04-201 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
70154 188/1944 04-202 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 09 July 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 09 July 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
70099 185/0255 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
70101 185/0658 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
70134 187/1152 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
70154 188/2004 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 09 July 2004
02 July 2004 - GMT Day 184 Beta = 66.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 +00.0/-33.6 +3.23/+2.29 +18.8/-12.8 27.0 1.00 15.6
3 +11.2/-33.6 +2.60/+1.66 +18.0/-12.8 27.0 1.01 15.4
09 July 2004 - GMT Day 191 Beta = 36.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.62 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-134.4 +5.15/+3.87 +26.4/-14.4 25.1 1.04 19.7
3 +22.4/-067.2 +3.87/+2.60 +25.6/-14.0 25.1 1.02 19.3
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 09 July 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-039 FOT 70177 185/1558 30832 Generic Late Acquisition
04-040 SOLSTICE 70140 187/2029 N/A SOLSTICE Microload Problem
04-039 FOT 70164 189/1143 30832 Generic Late Acquisition
04-041 FOT 70167 189/1621 36316 Late Acq. due to Guam equip
04-042 FOT 70177 190/0739 36321 Late Acq. Software prob @ WSGT
04-039 FOT 70183 190/1757 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-032 FOT See List June 04 30832 Generic Late Acquisition
04-034 PEM AXIS 70000 178/1357 N/A AXIS 1-2 OFF /ERRORS
04-035 PEM AXIS 70003 178/1729 N/A AXIS TURN ON/COP INITIAL
04-036 FOT 70003 178/1729 36240 TELEMETRY DROPOUT/WS WEATHER
04-037 FOT 70005 178/2057 36242 Late Acq - Bad Weather at Guam
04-038 FOT 70040 181/0448 36295 TLM Dropouts- Weather @ WSGT
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-039 FOT See List July 04 30832 Generic Late Acquisition
04-040 SOLSTICE 70140 187/2029 N/A SOLSTICE Microload Problem
04-041 FOT 70167 189/1621 36316 Late Acq. due to Guam equip
04-042 FOT 70177 190/0739 36321 Late Acq. Software prob @ WSGT